# ESIGMAHM: An Eccentric, Spinning inspiral-merger-ringdown waveform model with Higher Modes for the detection and characterization of binary black holes

### v1

@@ Abstract
With the updated sensitivity of terrestrial gravitational-wave (GW) detectors in their fourth observing run, we expect a high rate of detection of binary black hole mergers. With this comes the hope that we will detect rarer classes of merger events. Compact binaries formed dynamically in dense stellar environments are likely to be detected with a residual eccentricity as they enter the sensitivity band of ground-based GW detectors. In this paper, we present a time-domain inspiral-merger-ringdowm (IMR) waveform model ESIGMAHM constructed within a framework we named ESIGMA for coalescing binaries of spinning black holes on moderately eccentric orbits [1]. We now include the effect of black hole spins on the dynamics of eccentric binaries, as well as model sub-dominant waveform harmonics emitted by them. The inspiral evolution is described by a consistent combination of latest results from post-Newtonian theory, self-force, and black hole perturbation theory. We assume that these moderately eccentric binaries radiate away most of their orbital eccentricity before merger, and seamlessly connect the eccentric inspiral with a numerical relativity based surrogate waveform model for mergers of spinning binaries on quasi-circular orbits. We present two variants of ESIGMAHM: the inspiral-only version is named InspiralESIGMAHM, while the full IMR one is termed IMRESIGMAHM; or InspiralESIGMA and IMRESIGMA when only dominant modes are used. We validate ESIGMAHM against eccentric Numerical Relativity simulations, and also against contemporary effective-one-body and phenomenological models in the quasi-circular limit. We find that ESIGMAHM achieves match values greater than 99% for quasi-circular spin-aligned binaries with mass ratios up to 8, and above 97% for non-spinning and spinning eccentric systems with small or positively aligned spins. Using IMRESIGMA, we quantify the impact of orbital eccentricity on GW signals, showing that next-generation detectors can detect eccentric sources up to 10% louder than quasi-circular ones. We also show that current templated LIGO-Virgo-KAGRA searches will lose more than 10% of optimal SNR for about 20% of all eccentric sources by using only quasi-circular waveform templates. The same will result in a 25% loss in detection rate for eccentric sources with mass ratios $m_{1}/m_{2}\geq 4$. Our results highlight the need for including eccentricity and higher-order modes in GW source models and searches for asymmetric eccentric BBH signals.
@@ I Introduction
The detection of gravitational waves (GWs) by LIGO and Virgo detectors [2, 3] has opened a new window to probe the astrophysics of binary compact objects. By the end of the third observing run (O3), the ground-based network of detectors comprising of LIGO [4] and Virgo [5] detectors had detected nearly 100 compact binary coalescence (CBC) events [6, 7, 8, 9]. Compact binaries of stellar-mass black holes (BHs) and neutron stars (NSs) are prime targets for the existing and upcoming ground-based detectors such as LIGO A+ [10, 11], Voyager [11], Cosmic Explorer (CE) [12], and Einstein Telescope (ET) [13]. A majority of these binaries are expected to evolve from binary stars in galactic fields [14]. Their orbits are expected to circularize before reaching the frequency band of ground-based detectors by radiating eccentricity away during the long inspiral [15, 16]. However, for a sub-population of compact binaries formed via dynamical interaction in dense stellar environments [17, 18, 19, 20, 21] or through Kozai-Lidov processes [22, 23], this will not be the case. Such sources can have measurable residual eccentricities as they enter the sensitivity bands of ground-based detectors [6, 24, 25, 7, 8, 26, 27, 28, 29, 30, 31]. Indeed, recent follow-up analyses of GW190521 [32] claim this event to be consistent with a binary black hole (BBH) system with eccentricity ranging from $\sim 0.1-0.7$ [33, 34]. Carefully measuring the imprints of eccentricity in GW signals from CBCs will allow detailed studies of binary formation and evolution channels, as well as supporting astrophysical processes in globular clusters and galactic nuclei, which otherwise could remain inaccessible [18, 35, 36, 37, 38, 39, 40, 41, 42, 30, 43, 44, 45]. As the sensitivities of the current network of detectors [46] improve at lower frequencies, we expect to observe an ever-increasing number of GW events, improving the possibility of detecting an eccentric CBC population [47, 48].
@@
Detection of GW signals from CBCs in noise-dominated interferometric data involves matched filtering [49, 50, 51], where pre-computed waveforms (called templates) for the expected sources are cross-correlated with noisy detector output [52]. This matched-filtering technique is very sensitive to filter templates. Extracting information optimally from GW signals therefore requires highly accurate templates [53, 54, 55, 56, 57, 58] that include all physical phenomenology we expect our sources to exhibit [59]. The presence of eccentricity alters the length of the signal when compared to a circular counterpart with exact same parameters (see Fig. 9 of Ref. [60]). It also introduces amplitude and phase modulations at orbital timescale [61]. Spins of binary constituents also modify the size and shape of the waveform. For example, the length of the gravitational waveform for an (anti-) aligned spin BBH can be significantly (shorter) longer compared to a non-spinning one [62] (see Fig. 3-6 of Ref. [63]). Thus the combined effect of component spins and orbital eccentricity could be moderately degenerate [59]. Further, in order to describe binaries with asymmetric mass components, sub-dominant higher-order modes (HMs) play a crucial role. The inclusion of HMs introduces additional dependencies between binary parameters, which are useful in resolving degeneracies between different binary parameters such as luminosity distance and orbital inclination [64, 65, 66, 67]. HMs are also key to measure recoils due to BBH kicks, in particular its direction [68, 69, 70, 71], as well as for robust tests of General Relativity [72, 73].
@@
Contemporary GW data analyses therefore predominantly rely on waveform models that arise out of an amalgamation of multiple analytical approaches to the inspiraling two-body problem in GR, calibrated to NR simulations of the late stages of binary mergers [93]. Most of these models in literature have been traditionally focused on quasi-circular spinning compact binaries [94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115]. In the last few years, progress in constructing waveform models including eccentricity effects has also been made. A handful of eccentric inspiral-only [116, 117, 118, 100, 118, 119, 120, 121, 104] and inspiral-merger-ringdown (IMR) models [122, 123, 124, 1, 60, 125] have become available. We highlight some recent ones. The first, ENIGMA [122, 1], was aimed at moderate initial eccentricities (up to $\sim 0.4$), and combined an eccentric inspiral description with a quasi-circular merger model [122]. This paradigm was subsequently adopted by many modeling efforts. SEOBNRE modified an aligned-spin quasi-circular EOB waveform model (SEOBNRv1) to include effects of eccentricity in the radiative dynamics and waveform multipoles [126, 63]. Similarly, Ref. [125] modified a different aligned-spin EOB waveform model (TEOBResumS_SM[127, 128] to include effects of eccentricity. This model was further generalized by replacing the base quasi-circular model with one capable of describing generic-orbits [129]. There are other progresses in eccentric waveform modeling [130, 131, 132, 133] in surrogate models [92] and in EOB framework [134]. In addition to these models, Ref. [135] recently developed a method to add eccentricity induced modulations to existing quasi-circular BBH waveforms. The space of eccentric spinning IMR models is still sparsely populated in literature. And their accuracy has not been as rigorously tested as it has been for quasi-circular waveform models over the past decades [57, 56, 136].
@@
In this paper, we present a full IMR gravitational waveform model ESIGMAHM (Eccentric Spinning Inspiral with Generalized Merger Approximant and Higher-order Modes). It is a time-domain aligned-spin eccentric waveform model that includes higher gravitational wave harmonics, and is a significant extension of the ENIGMA framework [1]. ESIGMAHM carefully combines results from PN theory and self-force program to describe the low-frequency, weak field inspiral dynamics and GW emission. The high-frequency, strong field merger-ringdown (MR) uses a NR based surrogate model. The fundamental assumption we invoke remains that by the time the evolution of the binary is transitioning from inspiral to plunge and merger, it has lost most of its eccentricity via GW emission. This allows us to attach quasi-circular MR to an eccentric inspiral. The inspiral segment of ESIGMAHM has the latest spinning eccentric PN corrections reported in literature, and we use NRSur7dq4 [91] as our MR model, which is nearly as accurate as NR simulations themselves in describing quasi-circular BBH mergers. We present two versions of ESIGMAHM: the inspiral-only version is named InspiralESIGMAHM, while the full IMR one is termed IMRESIGMAHM; or InspiralESIGMA and IMRESIGMA when only dominant $\ell=|m|=2$ modes are chosen. Note that, we use IMRESIGMA and IMRESIGMAHM interchangeably in the rest of the paper to refer to our IMR model and clarify the inclusion of HMs wherever they’re used.
@@ II Executive Summary
The inaugural work [122] in the ENIGMA series introduced a non-spinning, moderately eccentric, dominant mode, time-domain, IMR waveform model. While the inspiral segment was constructed using a consistent combination of post-Newtonian, self-force and black hole perturbation theory results, the MR part was based on an implicit-rotating-source (IRS) model based on Ref. [139]. It reproduced the accurate dynamics of moderately eccentric BBH mergers with mass-ratios $q\in\{1,2\}$ from inspiral through merger. In the non-spinning limit, this waveform agreed well with an IMR effective-one-body (EOB) model  [63, 140].
@@
The second paper [1] introduced a Gaussian-process interpolated MR model named as Eccentric Non-spinning Inspiral Gaussian Merger Approximant (hence ENIGMA). The inspiral segment was improved by including the quasi-circular corrections in energy flux from black hole perturbation theory (BHPT) up to 6PN order. ENIGMA was validated using a set of Einstein Toolkit eccentric NR waveforms, with mass-ratios between $1\leq q\leq 5.5$, and eccentricities $e_{0}\leq 0.2$ ten orbits before the merger. ENIGMA reproduced the EOB model, SEOBNRv4 [109] in the non-spinning, quasi-circular limit with a match greater than $99\%$.
@@
Further, Ref. [60] upgraded the model by improving how the inspiral and MR portions are stitched together. This improved version reproduced the physics of quasi-circular mergers with very high accuracy over the applicable parameter space of ENIGMA.
@@
This paper extends these works [122, 1, 60] and introduces our new waveform model ESIGMAHM: Eccentric Spinning Inspiral Generalized Merger Approximant with Higher-order Modes, which incorporates component spin effects and includes sub-dominant waveform multipoles. ESIGMAHM includes recently computed high-order PN corrections to radiation reaction and gravitational waveform for compact binaries on eccentric orbits with spinning components [107] in the inspiral piece. It employs a recent NR surrogate model NRSur7dq4 [91] for quasi-circular spinning binaries as the MR approximant. This construction is valid for eccentric CBC sources that radiate away eccentricity by the end of inspiral to negligible levels. In rest of this paper we will refer to two variants of our model: InspiralESIGMAHM will refer to an inspiral-only version with waveform multipoles included up to $\ell=8$; and IMRESIGMAHM, which is the full IMR version that will include modes $(2,\pm 1),(3,\pm 3),(3,\pm 2),(4,\pm 4),(4,\pm 3)$ apart from dominant $(2,\pm 2)$ modes. IMRESIGMA can be used for the detection and follow-up of binaries with moderate eccentricities $e_{0}\leq 0.4$ [141, 107], mass-ratios up to $q\leq 6$ and positive-aligned or small anti-aligned component spins.
@@
We investigated the impact of orbital eccentricity on (GW) signals using the IMRESIGMA model. Our focus is on both current and third-generation ground-based GW observatories. We show that eccentric binaries exhibit an increased inspiral rate, leading to a significantly shorter time to merger compared to quasi-circular cases. Specifically, we show that the difference in the number of GW cycles increases very rapidly with mass ratio and initial eccentricity. Additionally, we analyze the optimal signal-to-noise ratio (SNR) for eccentric binaries, revealing that next-generation detectors like the Einstein Telescope and Cosmic Explorer can detect eccentric sources up to 10% louder than quasi-circular ones. We also show that current matched-filtering searches may miss 10% or more of optimal SNR for eccentric sources, and lead to a $25\%$ reduction in the detection rates of those with mass ratios above $4$. Note that the SNR losses we find due to the omission of HMs for eccentric systems are qualitatively consistent with the impact such omission has in non-eccentric cases [142, 143, 144, 145]. These results emphasize the importance of including orbital eccentricity effects and higher-order modes in GW waveform models as well as using such models in GW searches for BBHs.

### v2

@@ Abstract
With the updated sensitivity of terrestrial gravitational-wave (GW) detectors in their fourth observing run, we expect a high rate of detection of binary black hole mergers. With this comes the hope that we will detect rarer classes of merger events. Compact binaries formed dynamically in dense stellar environments are likely to be detected with a residual eccentricity as they enter the sensitivity band of ground-based GW detectors. In this paper, we present a time-domain inspiral-merger-ringdowm (IMR) waveform model ESIGMAHM constructed within a framework we named ESIGMA for coalescing binaries of spinning black holes on moderately eccentric orbits [1]. We now include the effect of black hole spins on the dynamics of eccentric binaries, as well as model sub-dominant waveform harmonics emitted by them. The inspiral evolution is described by a consistent combination of latest results from post-Newtonian theory, self-force, and black hole perturbation theory. We assume that these moderately eccentric binaries radiate away most of their orbital eccentricity before merger, and seamlessly connect the eccentric inspiral with a numerical relativity based surrogate waveform model NRSur7dq4 for mergers of spinning binaries on quasi-circular orbits. We present two variants of ESIGMAHM: the inspiral-only version is named InspiralESIGMAHM, while the full IMR one is termed IMRESIGMAHM; or InspiralESIGMA and IMRESIGMA when only dominant modes are used. We validate ESIGMAHM against eccentric Numerical Relativity simulations, and also against contemporary effective-one-body and phenomenological models in the quasi-circular limit. We find that ESIGMAHM achieves match values greater than 99% for quasi-circular spin-aligned binaries with mass ratios up to 8, and above 97% for non-spinning and spinning eccentric systems with small or positively aligned spins. Using IMRESIGMA, we quantify the impact of orbital eccentricity on GW signals, showing that next-generation detectors can detect eccentric sources up to 10% louder than quasi-circular ones. We also show that current templated LIGO-Virgo-KAGRA searches will lose more than 10% of optimal SNR for about 20% of all eccentric sources by using only quasi-circular waveform templates. The same will result in a 25% loss in detection rate for eccentric sources with mass ratios $m_{1}/m_{2}\geq 4$. Our results highlight the need for including eccentricity and higher-order modes in GW source models and searches for asymmetric eccentric BBH signals.
@@ I Introduction
The detection of gravitational waves (GWs) by LIGO and Virgo detectors [2, 3] has opened a new window to probe the astrophysics of binary compact objects. By the end of the third observing run (O3), the ground-based network of detectors comprising of LIGO [4] and Virgo [5] detectors had detected more than 90 compact binary coalescence (CBC) events [6, 7, 8, 9]. Compact binaries of stellar-mass black holes (BHs) and neutron stars (NSs) are prime targets for the existing and upcoming ground-based detectors such as LIGO A+ [10, 11], Voyager [11], Cosmic Explorer (CE) [12], and Einstein Telescope (ET) [13]. A majority of these binaries are expected to evolve from binary stars in galactic fields [14]. Their orbits are expected to circularize before reaching the frequency band of ground-based detectors by radiating eccentricity away during the long inspiral [15, 16]. However, for a sub-population of compact binaries formed via dynamical interaction in dense stellar environments [17, 18, 19, 20, 21] or through Kozai-Lidov processes [22, 23], this will not be the case. Such sources can have measurable residual eccentricities as they enter the sensitivity bands of ground-based detectors [6, 24, 25, 7, 8, 26, 27, 28, 29, 30, 31]. Indeed, recent follow-up analyses of GW190521 [32, 33, 34] claim this event to be consistent with a binary black hole (BBH) system with eccentricity ranging from $\sim 0.1-0.7$ [35, 36, 37]. Carefully measuring the imprints of eccentricity in GW signals from CBCs will allow detailed studies of binary formation and evolution channels, as well as supporting astrophysical processes in globular clusters and galactic nuclei, which otherwise could remain inaccessible [18, 38, 39, 40, 41, 42, 43, 44, 45, 30, 46, 47, 48]. As the sensitivities of the current network of detectors [49] improve at lower frequencies, we expect to observe an ever-increasing number of GW events, improving the possibility of detecting an eccentric CBC population [50, 51].
@@
Detection of GW signals from CBCs in noise-dominated interferometric data involves matched filtering [52, 53, 54], where pre-computed waveforms (called templates) for the expected sources are cross-correlated with noisy detector output [55]. This matched-filtering technique is very sensitive to filter templates. Extracting information optimally from GW signals therefore requires highly accurate templates [56, 57, 58, 59, 60, 61] that include all physical phenomenology we expect our sources to exhibit [37]. The presence of eccentricity alters the length of the signal when compared to a circular counterpart with exact same parameters (see Fig. 9 of Ref. [62]). It also introduces amplitude and phase modulations at orbital timescale [63]. Spins of binary constituents also modify the size and shape of the waveform. For example, the length of the gravitational waveform for an (anti-) aligned spin BBH can be significantly (shorter) longer compared to a non-spinning one [64] (see Fig. 3-6 of Ref. [65]). Thus the combined effect of component spins and orbital eccentricity could be moderately degenerate [37]. Further, in order to describe binaries with asymmetric mass components, subdominant higher-order modes (HMs) play a crucial role. The inclusion of HMs introduces additional dependencies between binary parameters, which are useful in resolving degeneracies between different binary parameters such as luminosity distance and orbital inclination [66, 67, 68, 69]. HMs are also key to measure recoils due to BBH kicks, in particular its direction [70, 71, 72, 73], as well as for robust tests of General Relativity [74, 75].
@@
Contemporary GW data analyses therefore predominantly rely on waveform models that arise out of an amalgamation of multiple analytical approaches to the coalescing two-body problem in GR, calibrated to NR simulations of the late stages of binary mergers [94]. Most of these models in literature have been traditionally focused on quasi-circular spinning compact binaries [95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123]. In the last few years, especially since the discovery of GW150914, progress in constructing waveform models including eccentricity effects has also been made [124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 130, 146, 147, 148, 134, 149, 150, 1, 151, 1, 152, 65, 62, 153, 101, 104, 154, 155, 156, 157, 158, 159, 93, 160, 161, 142]. A series of eccentric inspiral-only [143, 144, 145, 130, 146, 147, 148, 134] and inspiral-merger-ringdown (IMR) models [149, 150, 1, 151, 1, 152, 65, 62, 153, 101, 104, 154, 155, 156, 157, 158, 159, 93, 160, 161, 142, 141, 140, 139] have become available. We highlight some recent ones. The first IMR model, ENIGMA [149, 1], was aimed at moderate initial eccentricities (up to $\sim 0.4$), and combined an eccentric inspiral description with a quasi-circular merger model [149]. This paradigm was subsequently extended by many modeling efforts. SEOBNRE modified an aligned-spin quasi-circular EOB waveform model (SEOBNRv1) to include effects of eccentricity in the radiative dynamics and waveform multipoles [152, 65]. Similarly, Ref. [153] modified a different aligned-spin EOB waveform model (TEOBResumS_SM[101, 104] to include effects of eccentricity. This model was further generalized by replacing the base quasi-circular model with one capable of describing generic-orbits [154]. All of these models extend upon the paradigm of ENIGMA by using an eccentric description all the way up to the binary merger. There are other progresses in eccentric waveform modeling [156, 157, 158, 159] in surrogate models [93, 160] and in the EOB framework [155]. In addition to these models, Ref. [162, 163, 164, 165], recently developed a method to add eccentricity induced modulations to existing quasi-circular BBH waveforms. The space of the eccentric spinning IMR models is sparsely populated in the literature though, especially since long eccentric simulations are the ones required for meaningful comparison with PN/BHP/self-force solutions [166, 167]. And their accuracy has not been as rigorously tested as it has been for quasi-circular waveform models over the past decades [60, 59, 168].
@@
In this paper, we present a full IMR gravitational waveform model ESIGMAHM (Eccentric Spinning Inspiral with Generalized Merger Approximant and Higher-order Modes). It is a time-domain aligned-spin eccentric waveform model that includes higher gravitational wave harmonics, and is a significant extension of the ENIGMA framework [1]. ESIGMAHM carefully combines results from PN theory and self-force program to describe the low-frequency, weak field inspiral dynamics and GW emission. The high-frequency, strong field plunge-merger-ringdown (PMR) uses a NR based surrogate model. The fundamental assumption we invoke remains that by the time the evolution of the binary is transitioning from inspiral to plunge and merger, it has lost most of its eccentricity via GW emission. This allows us to attach quasi-circular PMR to an eccentric inspiral. The inspiral segment of ESIGMAHM has the latest spinning eccentric PN corrections reported in literature, and we use NRSur7dq4 [92] as our PMR model, which is nearly as accurate as NR simulations themselves in describing quasi-circular BBH mergers. We present two versions of ESIGMAHM: the inspiral-only version is named InspiralESIGMAHM, while the full IMR one is termed IMRESIGMAHM; or InspiralESIGMA and IMRESIGMA when only dominant $\ell=|m|=2$ modes are chosen. Note that, we use IMRESIGMA and IMRESIGMAHM interchangeably in the rest of the paper to refer to our IMR model and clarify the inclusion of HMs wherever they’re used.
@@ II Executive Summary
The inaugural work [149] in the ENIGMA series introduced a non-spinning, moderately eccentric, dominant mode, time-domain, IMR waveform model. While the inspiral segment was constructed using a consistent combination of post-Newtonian, self-force and black hole perturbation theory results, the PMR part was based on an implicit-rotating-source (IRS) model based on Ref. [171, 172]. It reproduced the accurate dynamics of moderately eccentric BBH mergers with mass-ratios $q\in\{1,2\}$ from inspiral through merger. In the non-spinning limit, this waveform agreed well with an IMR effective-one-body (EOB) model  [65, 173].
@@
The second paper [1] introduced a Gaussian-process interpolated IMR model named as Eccentric Non-spinning Inspiral Gaussian Merger Approximant (hence ENIGMA). The inspiral segment was improved by including the quasi-circular corrections in energy flux from black hole perturbation theory (BHPT) up to 4PN order. ENIGMA was validated using a set of Einstein Toolkit eccentric NR waveforms, with mass-ratios between $1\leq q\leq 5.5$, and eccentricities $e_{0}\leq 0.2$ ten orbits before the merger. ENIGMA reproduced the EOB model, SEOBNRv4 [100] in the non-spinning, quasi-circular limit with a match greater than $99\%$.
@@
Further, Ref. [62] upgraded the model by improving how the inspiral and PMR portions are stitched together. This improved version reproduced the physics of quasi-circular mergers with very high accuracy over the applicable parameter space of ENIGMA.
@@
This paper extends these works [149, 1, 62] and introduces our new waveform model ESIGMAHM: Eccentric Spinning Inspiral Generalized Merger Approximant with Higher-order Modes, which incorporates component spin effects and includes sub-dominant waveform multipoles. ESIGMAHM includes recently computed high-order PN corrections to radiation reaction and gravitational waveform for compact binaries on eccentric orbits with spinning components [138] in the inspiral piece. It employs a recent NR surrogate model NRSur7dq4 [92] for quasi-circular spinning binaries as the PMR approximant. This construction is valid for eccentric CBC sources that radiate away eccentricity by the end of inspiral to negligible levels. In rest of this paper we will refer to two variants of our model: InspiralESIGMAHM will refer to an inspiral-only version with waveform multipoles included up to $\ell=8$; and IMRESIGMAHM, which is the full IMR version that will include modes $(2,\pm 1),(3,\pm 3),(3,\pm 2),(4,\pm 4),(4,\pm 3)$ apart from dominant $(2,\pm 2)$ modes. IMRESIGMA can be used for the detection and follow-up of binaries with moderate eccentricities $e_{0}\leq 0.4$ [174, 138], mass-ratios up to $q\leq 6$ and positive-aligned or small anti-aligned component spins.
@@
We investigated the impact of orbital eccentricity on (GW) signals using the IMRESIGMA model. Our focus is on both current and third-generation ground-based GW observatories. We show that eccentric binaries exhibit an increased inspiral rate, leading to a significantly shorter time to merger compared to quasi-circular cases. Specifically, we show that the difference in the number of GW cycles increases very rapidly with mass ratio and initial eccentricity. Additionally, we analyze the optimal signal-to-noise ratio (SNR) for eccentric binaries, revealing that next-generation detectors like the Einstein Telescope and Cosmic Explorer can detect eccentric sources up to 10% louder than quasi-circular ones. We also show that current matched-filtering searches may miss 10% or more of optimal SNR for eccentric sources, and lead to a $25\%$ reduction in the detection rates of those with mass ratios above $4$. Note that the SNR losses we find due to the omission of HMs for eccentric systems are qualitatively consistent with the impact such omission has in non-eccentric cases [175, 176, 177, 178]. These results emphasize the importance of including orbital eccentricity effects and higher-order modes in GW waveform models as well as using such models in GW searches for BBHs. Further, a long term goal of the ESIGMA framework would be to iteratively develop better waveform models for the eccentric inspiral/plunge and ultimately use eccentric NR surrogates to represent the plunge-merger-ringdown.